Suppose the current annual salary of all teachers in the Philippines have a normal distribution with a mean of 95,000 pesos and a standard deviation of 20,000 pesos.

 

a.  Find the probability that the annual salary of a randomly selected teacher would be between 56,000 and 76,000.

 

b.  Find the probability that the annual salary of a randomly selected teacher would be at least 50,000 pesos.

 

c.  Find the probability that the annual salary of a randomly selected teacher would be 126,000 pesos.


1
Expert's answer
2022-05-02T09:13:10-0400

Let "X=" the current annual salary: "X\sim N(\mu, \sigma^2)."

Given "\mu=95000, \sigma=20000."

a.

"P(56000<X<76000)"

"=P(Z<\dfrac{76000-95000}{20000})"

"-P(Z\le\dfrac{56000-95000}{20000})\approx0.1455"

b.

"P(X\ge50000)=1-P(Z<\dfrac{50000-95000}{20000})"

"\approx0.9878"

c.


"P(X=126000)=0"